# Revisiting Randomization in Greedy Model Search

### v1

@@ 4.2 Comparisons with Elastic Net
It is instructive to compare the behavior of RGS with that of elastic net regression. Introduced in [62], elastic net is a popular method that combines the $\ell_{1}$ and $\ell_{2}$ penalties of lasso and ridge regression, respectively. Specifically, it solves the convex optimization problem
$\hat{\boldsymbol{\beta}}\phantom{}^{\text{EN}}=\operatorname*{arg\,min}_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\Big{\{}\|\mathbf{y}-\mathbf{X}\boldsymbol{\beta}\|^{2}+\lambda_{1}\|\boldsymbol{\beta}\|_{\ell_{1}}+\lambda_{2}\|\boldsymbol{\beta}\|^{2}_{\ell_{2}}\Big{\}},$ (16)
where $\lambda_{1}$ and $\lambda_{2}$ are non-negative tuning hyperparameters. Under Assumption 3.1, the elastic net estimator takes a particularly simple form. As shown in [62], the elastic net coefficient for $x_{j,p}$ can be expressed as
$\hat{\beta}\phantom{}^{\text{EN}}_{j,p}=\frac{(|\hat{\beta}_{j,p}|-n\lambda_{1}/2)_{+}}{1+n\lambda_{2}}\text{sgn}(\hat{\beta}_{j,p}),$ (17)
where $\hat{\beta}_{j,p}$ is the ordered OLS coefficient from Definition 3.1. For comparison, using the representation (6) and approximation (7), the RGS coefficient for $x_{j,p}$ has the form:
$\hat{\beta}\phantom{}^{\text{RGS}}_{j,p}\approx\frac{\hat{\beta}_{j,p}}{1+\big{(}1-\frac{m}{p}\big{)}^{k-j+1/2}}.$ (18)
@@
Elastic net was originally designed to relax lasso’s preference for highly sparse solutions, which was observed to create instability when features are highly correlated. From (17), we see that it has the same effect even in an orthogonal setting. Since all non-zero coefficients are shrunk by a constant factor of $1/(1+n\lambda_{2})$, the parameter budget of the model can be spread out across more coefficients. While this is somewhat similar in spirit to the role of the feature subsampling in RGS, the family of solutions attainable using both strategies is different, thereby yielding different predictive performances on any given dataset.
4.3 Revisiting Best Subset vs Lasso
@@
In the sparse linear regression literature, Hastie, Tibshirani, and Tibshirani [32] compared BSS and lasso, showing via numerical experiments that in high-SNR settings, BSS (and FS) seem to have higher accuracy than lasso. In contrast, in low-SNR regimes, they showed that lasso outperforms BSS. [^10] See [29] for some recent theoretical work supporting this observation. They explain this by observing that “different procedures bring us from the high-bias to the high-variance ends of the tradeoff along different model paths”, with the solution path of neither algorithm completely dominating the other. Our weight calculations, together with our earlier theoretical results, adds to their discussion by showing that the BSS solution path is not Pareto optimal. The RGS weights also offer an alternative solution path to those traversed by lasso and elastic net, thereby revealing how these may potentially have limitations.
@@ 4.4 Shrinkage
While decreasing the feature subsampling parameter $m$ does not necessarily shrink each coefficient in the RGS solution (since some weights increase while others decrease), it does have the effect of shrinking the fitted values towards zero. We state this fact in terms of the following theorem, whose proof relies on establishing that the RGS weight sequences define a majorization hierarchy parameterized by $m$. Evidently, the largest shrinkage occurs when $m=1$, in which case $\|\hat{\mathbf{f}}_{k,1,p}\|=\sqrt{k/p}\|\hat{\boldsymbol{\beta}}\|_{\ell_{2}}$.
Theorem 4.3.
Grant Assumption 3.1 and let $m<m^{\prime}$. Then, for any $k,p\geq 1$, we have
$\|\hat{\mathbf{f}}_{k,m,p}\|\leq\|\hat{\mathbf{f}}_{k,m^{\prime},p}\|.$ (19)
5 Numerical Experiments
In this section, we describe the results of several numerical experiments that illustrate the predictive effectiveness of RGS vis-à-vis other sparse regression methods. Furthermore, our experiments provide evidence that our theoretical results generalize beyond the idealized assumptions, such as orthogonal features, required in Section 3. We present only the key settings and findings, with additional experiments found in the supplement.
5.1 Experimental Design
5.1.1 Data Generating Process
We generate i.i.d. training data $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, where $\mathbf{x}_{i}\in\mathbb{R}^{p}$ has the distribution $\mathbf{x}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma})$, $y_{i}=\boldsymbol{\beta}^{\top}\mathbf{x}_{i}+\varepsilon_{i}$, with $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, independent of $\mathbf{x}_{i}$. Following Hastie et al. [32], we investigate two covariance structures: (i) banded correlation, in which the $(i,j)$ entry of $\boldsymbol{\Sigma}$ satisfies $\Sigma_{ij}=\rho^{|i-j|}$, and (ii) block correlation, in which $\Sigma_{ii}=1$, $\Sigma_{ij}=0.25$ if features $i$ and $j$ lie in the same block, and $\Sigma_{ij}=0$ otherwise. We investigate two types of regression vectors: (i) exact sparsity, whereby $\beta_{i}=1$ for $1\leq i\leq s$ and $\beta_{i}=0$ otherwise, and (ii) inexact sparsity, whereby $\beta_{i}=1$ for $1\leq i\leq s$ and $\beta_{i}=(-1)^{i}\exp(-i/2)$ for $s<i\leq p$. Note that here we use a fixed ordering of the features instead of the data-dependent ordering used in our theoretical discussions. We fix $n=1000$, $s=10$, and vary $p\in\{100,200,400,500,800\}$, $\rho\in\{0.10,0.25,0.50,0.75\}$. Furthermore, we vary $\sigma^{2}$ so that the signal-to-noise ratio (SNR), defined as $\textnormal{SNR}\coloneqq\mathbb{E}[(\boldsymbol{\beta}^{\top}\mathbf{x})^{2}]/\sigma^{2}=\boldsymbol{\beta}^{\top}\boldsymbol{\Sigma}\boldsymbol{\beta}/\sigma^{2}$, lies in the grid $\{0.031,0.053,0.11,0.25,1.0\}$. These values correspond to a roughly geometric grid $\{0.03,0.05,0.1,0.2,0.5\}$ under the normalized scaling $\textnormal{SNR}/(\textnormal{SNR}+1)$. [^11] This transformed quantity is sometimes called the “perfect score” proportion of variance explained (PVE) [31, 1] and is also known as heritability in genomics [52].
5.1.2 Methods and Hyperparameters
We compare RGS against GS (FS) as well as two other randomized ensembling methods for GS: (i) bagging [7] and (ii) smearing [9]. We also compare against lasso regression [50] and elastic net [62], which are two of the most commonly used convex relaxation methods for sparse regression. We optimized the hyperparameters for all the methods based on 10-fold cross validation. We varied the hyperparameters $(k,m)$ for RGS by considering every $k\in[20]$ and $m\in\{\lfloor 2+(p-2)(1.5^{i}-1)/(1.5^{9}-1)\rfloor\}_{i=0}^{9}$. This formula for $m$ was chosen to consider values growing exponentially, with more values considered near $1$ than $p$, as standard literature on random forests suggest that lower values of $m$, particularly $m=1/3$, are closer to the optimum. The number of base learners in each ensembling method was set at $B=500$.
5.1.3 Evaluation Metrics
We evaluated the performance of each method according to two metrics: (i) the relative in-sample error (RISE), defined as $\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{\hat{\boldsymbol{\Sigma}}}^{2}/\sigma^{2}+1$, where $\hat{\boldsymbol{\Sigma}}$ is the sample covariance $\mathbf{X}^{\top}\mathbf{X}/n$, and (ii) the relative test error (RTE), defined as $\|\hat{\boldsymbol{\beta}}-\boldsymbol{\beta}\|_{\boldsymbol{\Sigma}}^{2}/\sigma^{2}+1$. The RISE can be shown to be equivalent to the value $\mathbb{E}[\|\hat{\mathbf{f}}-\mathbf{y}^{\prime}\|^{2}]/\sigma^{2}$, where $\hat{\mathbf{f}}$ is the vector of fitted values, and $y_{i}^{\prime}=f(\mathbf{x}_{i})+\varepsilon_{i}^{\prime}$, with $\varepsilon_{i}^{\prime}\sim\mathcal{N}(0,\sigma^{2})$ drawn independently of the training data. It hence measures the ratio of the in-sample error to the Bayes error. Similarly, the RTE can be shown to be equivalent to $\mathbb{E}_{(\mathbf{x},y)}[(\hat{\boldsymbol{\beta}}^{\top}\mathbf{x}-y)^{2}]/\sigma^{2}$, where $(\mathbf{x},y)$ is a fresh example drawn from the same distribution as the training set. This is hence the ratio of the test error to the Bayes error. We include both metrics because while the former is more relevant to fixed design linear regression, which is the framework used in our theoretical analysis, the latter is more relevant to modern machine learning. The reported values for both metrics were computed by averaging over 10 experimental replicates.
5.2 Results
In Figure 3, we show the RISE values for all methods under exact sparsity, banded correlation with $\rho=0.5$, and aspect ratios $p/n=0.1,0.8$, as SNR is varied over the full grid of values. All other results are deferred to Section B in the supplement. Similar to Hastie et al. [32], we observe that GS tends to outperform lasso at high SNR values, but suffers from poor performance at low SNR values. On other hand, RGS seems to uniformly outperform GS, and is competitive with both lasso and elastic net at low SNR values. Furthermore, it tends to outperform smearing and has similar performance to bagging, but with much higher computational efficiency, with RGS being about 5x faster than bagging per model fit on an AMD Ryzen 7 7800X3D CPU without parallelization. Finally, the improvement of RGS over other methods seems to be larger at a higher aspect ratio. These observations are relatively robust to the choice of aspect ratio, covariance structure, sparsity structure, and the choice of evaluation metric. In additional experiments, we also found our conclusions to hold even with non-Gaussian and heavy-tailed noise.
(a) $p/n=0.1$
(b) $p/n=0.8$
Figure 3: The relative in-sample error for various sparse regression methods across SNR $\in\{0.031,0.053,0.11,0.25,1.0\}$ for banded exact sparsity data. RGS together with bagged GS seem to uniformly outperform all other methods.
5.3 Bias-Variance Trade-Off
Towards validating Theorem 3.1 and Theorem 3.2 under non-orthogonal designs, we computed the degrees of freedom and training error for RGS for each fixed value of $k$, with $m$ selected via cross-validation, and compared them against the corresponding values for GS for the same value of $k$. We performed this for each simulation setting discussed in Section 5.1.1. The results for exact sparsity, banded correlated features with $\rho=0.5$, $\textnormal{SNR}=0.053,0.25,1.0$, and $p/n=0.1$ are shown in Figure 4. We observe that for $\textnormal{SNR}=0.053,0.25$, for each model size $k$, using RGS over GS decreases both the degrees of freedom and training error, thereby shifting the entire bias-variance tradeoff curve towards the bottom left. The gains are more significant when SNR is lower, which is in accordance with Figure 3. On the other hand, for high SNR (such as SNR $=1$), RGS tends to perform more poorly around $k=10$. This phenomenon is partly explained in Proposition 6.1 to come. Additional results for other settings are deferred to Section B in the supplement.
(a) SNR $=0.053$
(b) SNR $=0.25$
(c) SNR $=1.0$
Figure 4: Degrees of freedom (y-axis) vs. training MSE (x-axis) for RGS and GS with $p/n=0.1$, banded exact sparsity, as we vary $k\in[20]$, with the “true” value being 10. For each value of $k$, RGS decreases both the degrees of freedom as well as training MSE, thereby shifting the tradeoff curve towards the bottom left. The gains are more significant at low SNR. At high SNR, RGS and GS become similar.
6 Explaining the Role of Randomization in RFs
RFs [10] consist of ensembles of regression (or classification) trees, each of which recursively partitions the feature space, with each split (on a particular feature and at a particular threshold) selected by greedily optimizing a local performance criterion [11]. Each tree is randomized by bootstrapping the training dataset (bagging) and by randomly constraining the set of features that can be split on at each iteration, that is, by feature subsampling. Because of their resounding empirical success [15, 26, 44], there has been a significant amount of work on understanding the statistical and mathematical properties of RFs, such as on their consistency, convergence rates, and asymptotic normality, under various data generating scenarios [49, 41, 6, 35, 16, 2]. On the other hand, the exact role played by randomization in the effectiveness of RFs remains unclear, with there being much disagreement and debate [10, 31, 55, 42, 18, 37, 40, 48]. In this section, we summarize these developments and discuss how our results provide empirical and theoretical evidence for or against various arguments.
Breiman [10] originally explained the benefits of randomization as increasing diversity among the trees, thereby reducing the variance in their predictions. Interestingly, in his concluding remarks, he speculated that randomization may “act to reduce bias” but that “the mechanism for this is not obvious”. These two explanations have had divergent fates, with Hastie, Tibshirani, and Friedman [31] arguing in their influential textbook that “the improvements in prediction obtained by bagging or random forests are solely a result of variance reduction”.
In 2020, Mentch and Zhou elaborated on this hypothesis. They focused on the role played by feature subsampling and (i) found in their numerical experiments that decreasing the feature subsampling ratio $\gamma$ decreases the degrees of freedom in RFs. To further explore the generality of this phenomenon, they considered a version of RGS that also incorporates bagging, (ii) finding that the degrees of freedom in RGS also decreases with $\gamma$. They hence (iii) concluded that feature subsamping “serves as a form of implicit regularization” and that $\gamma$ “serves much the same purpose as the shrinkage penalty in explicitly regularized regression procedures like lasso and ridge regression”. Their claim inspired and was seemingly supported by a wave of theoretical work analyzing ensembles of (non-adaptive) OLS estimators randomized via feature subsampling, with the broad conclusion that the resulting estimator is mathematically equivalent to ridge regression [36, 47], in an asymptotic sense.
Our results suggest, contrary to this line of work, that the analogy between $\gamma$ and $\lambda$, the shrinkage penalty in ridge or lasso regression, is more nuanced in the context of greedy, adaptive methods like RGS (and RFs). Note that in both ridge and lasso regression, the expected training error (resp. degrees of freedom) is monotonically increasing (resp. monotonically decreasing) in $\lambda$ [63, 31], thereby giving an exact trade-off between the two quantities. However, this is not what happens with $\gamma$ in RGS. Our Theorem 3.2 shows that the training error for RGS is not monotonic in $\gamma$ and instead is minimized at an intermediate value. Meanwhile, although Theorem 3.2 in our paper gives a sufficient condition for which the degrees of freedom is monotonically decreasing in $\gamma$, which generally holds in lower SNR regimes and seems to corroborate observation (ii) above, it is easy to construct an example where this monotonicity fails in higher SNR regimes:
@@ 6 Explaining the Role of Randomization in RFs
Suppose that the regression vector $\boldsymbol{\beta}$ is exactly $s$-sparse (all nonzero coefficients are equal to 1) and that the noise variables are Gaussian. Then we have
$\lim\limits_{\begin{subarray}{c}m,\,p,\,n\to\infty,\\ m/p\to\gamma\end{subarray}}\frac{\operatorname{df}(\hat{f}_{k,m,p})-\operatorname{df}(\hat{f}_{k,p})}{\log p}\geq C,$ (20)
where $C$ is a constant depending only on $\gamma$ and $k$.
The breakdown in the analogy between $\gamma$ and $\lambda$ can also be seen in the fitted coefficients. While the ridge or lasso coefficients shrink toward zero as $\lambda$ increases, the weight formulas (Theorems 4.1 and 4.2) for RGS show that varying $\gamma$ simultaneously shrinks some coefficients while inflating others. In summary, the role of feature subsampling in RGS is altogether more complex than providing shrinkage. As mentioned in previous sections, the flexibility of being able to vary $\gamma$ has the effect of enlarging the model search space, which can lead to simultaneously better model fit and lower model complexity.
While our theory applies only to RGS, it coheres with the claims of a number of very recent works that randomization in RFs can reduce “bias” [18, 37, 40, 48]. While each of these works uses a different definition of bias, all definitions become equivalent to the classical definition if, similar to our analysis of RGS, we consider RFs in the limit of infinite trees, which exists by the law of large numbers. In this case, the bias of the RF estimator $\hat{f}$ is defined as $\mathbb{E}\big{[}\big{(}f(\mathbf{x}_{0})-\mathbb{E}\big{[}\hat{f}(\mathbf{x}_{0})\big{]}\big{)}^{2}\big{]}$, where the inner expectation is taken with respect to the training set and the outer expectation is taken with respect to a new data point $\mathbf{x}_{0}$.
In these works, we identify two mechanisms for bias reduction. The first mechanism is by enhancing the approximation power of RFs. As observed Curth et al. [18], the class of decision tree models that can be fitted on a fixed dataset is not convex, so by taking simple averages of these decision trees, we enlarge the class of possible models. [^12] To explain further, any decision tree model fitted on a fixed dataset is constrained to have leaf values that are equal to the average response in the training set over that leaf. In particular, they cannot have empty leaves. It is easy to see that this constraint makes this class of functions non-convex. Via numerical experiments, they showed that when varying $\gamma$ while fixing all other hyperparameters, minimum bias can be achieved at intermediate value of $\gamma$. Mei et al. [40] gave a theoretical argument to support this statement when the true regression function is linear. [^13] Their proof, however, relies on an oracle tree-growing process. Note that this argument is entirely analogous to our explanation for how randomization in RGS improves model fit (c.f., Remark Remark.) A second mechanism is by overcoming potential failures of greedy optimization, which is used by the base learners of RFs. Specifically, Liu and Mazumder [37] showed numerically how certain “hidden patterns” that would otherwise be missed can be detected upon injecting randomness into the tree-growing process, leading to smaller bias. While this mechanism has no analogy in our theoretical analysis of RGS (again c.f., the remark in Section 3.A.), it is easy to construct an example showing how it may arise in RGS as well (c.f., Example A.1 in the supplement).

### v3

@@ 5.1 Comparison with elastic net
It is instructive to compare EFS($k,m$) with elastic net regression (zou2005regularization), which combines the $\ell_{1}$ and $\ell_{2}$ penalties of lasso and ridge,
$\hat{\boldsymbol{\beta}}^{\text{EN}}=\operatorname*{arg\,min}_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\Big\{\|\mathbf{y}-\mathbf{X}\boldsymbol{\beta}\|^{2}+\lambda_{1}\|\boldsymbol{\beta}\|_{\ell_{1}}+\lambda_{2}\|\boldsymbol{\beta}\|^{2}_{\ell_{2}}\Big\}.$
Under Assumption 3.1, elastic net takes a simple form (zou2005regularization),
$\hat{\beta}^{\text{EN}}_{j,p}=\frac{(|\hat{\beta}_{j,p}|-\lambda_{1}/2)_{+}}{1+\lambda_{2}}\operatorname{sgn}(\hat{\beta}_{j,p}).$ (15)
For comparison, Theorem 3.2 implies the EFS($k,m$) coefficient is approximately
$\hat{\beta}^{\text{EFS}}_{j,p}\approx\frac{\hat{\beta}_{j,p}}{1+\big(1-\frac{m}{p}\big)^{k-j+1/2}}.$
@@
Elastic net was originally designed to mitigate lasso’s tendency toward highly sparse solutions, which causes instability when features are correlated. From (15), we see that even in the orthogonal setting, elastic net spreads the parameter budget across more coefficients by shrinking all nonzero coefficients by the constant factor $1/(1+\lambda_{2})$. This is similar in spirit to feature subsampling in EFS($k,m$), but the families of attainable solutions differ, yielding different predictive performances on any given dataset.
5.2 Revisiting best subset vs. lasso
@@
hastie2020extended compared BSS and lasso, showing that BSS (and FS($k$)) achieves higher accuracy than lasso in high-SNR settings, and that lasso dominates in low-SNR regimes. They explain this by observing that “different procedures bring us from the high-bias to the high-variance ends of the tradeoff along different model paths,” with neither algorithm’s solution path completely dominating the other. See also ghosh2025signal for recent theoretical support.
Our results add to this discussion by showing that the BSS solution path is not Pareto optimal. EFS can simultaneously reduce both bias and variance. The ensemble weights trace yet another solution path, distinct from those of lasso and elastic net.
@@ 5.3 Shrinkage
Decreasing $m$ does not shrink every coefficient (some weights increase and others decrease), but it does shrink the fitted values toward zero.
@@
Grant Assumption 3.1. For $m<m^{\prime}$ and any $k,p\geq 1$,
$\|\hat{\mathbf{f}}_{k,m,p}\|\leq\|\hat{\mathbf{f}}_{k,m^{\prime},p}\|.$
Proof sketch.
From (6), it suffices to show $\sum_{j=1}^{p}\hat{\beta}_{j,p}^{2}(w_{j}^{k,m,p})^{2}\leq\sum_{j=1}^{p}\hat{\beta}_{j,p}^{2}(w_{j}^{k,m^{\prime},p})^{2}$. This follows from the majorization established in Theorem 4.2 together with the Schur-convexity of $g(w_{1},\ldots,w_{p})=\sum_{j}\beta_{j}^{2}w_{j}^{2}$. ∎
The extremes are as follows. Maximal shrinkage occurs at $m=1$, where $\|\hat{\mathbf{f}}_{k,1,p}\|=(k/p)\|\hat{\boldsymbol{\beta}}\|_{\ell_{2}}$. No shrinkage occurs at $m=p$, where EFS($k,m$) reduces to FS($k$).
6 Another mechanism for training error reduction
Theorem 4.1 establishes training error reduction via the “enlarged model class” mechanism of curth2024random. The orthogonality assumption makes FS($k$) equivalent to BSS, so randomization helps only by searching over a richer family of coefficient profiles. However, liu2025randomization identified a second mechanism. Feature subsampling can help individual base estimators escape local optima that greedy optimization would otherwise get trapped in.
We now construct an example demonstrating this second mechanism. In the orthogonal setting, FS($k$) always selects the $k$ features most correlated with the response, which means that individual EFS base estimators cannot outperform it. We therefore introduce correlation in the form of a “spurious” feature $x_{p,p}$ that is strongly correlated with the response but not part of the optimal representation. Greedy FS selects this spurious feature, but EFS base estimators may avoid it if $x_{p,p}$ is excluded from the first candidate set.
Proposition 6.1 (Escaping local minima).
Consider a design matrix $\mathbf{X}$ with $\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot j,p}\rangle=\delta_{ij}$ for $1\leq i,j\leq{p-1}$, $\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot p,p}\rangle=3^{-1/2}$ for $i\in\{1,2\}$, and $\langle\mathbf{x}_{\cdot i,p},\mathbf{x}_{\cdot p,p}\rangle=0$ otherwise. Assume the noiseless model $y=\beta(x_{1,p}+x_{2,p})+\zeta\sum_{i=3}^{p-1}x_{i,p}$ with $6^{-1/2}\beta>\zeta>0$. Let $\check{f}_{k,m,p}$ denote a single EFS base estimator. There exists $C>0$ depending only on $\beta,\xi$ such that for $k,p\geq C$,
$\mathbb{E}\big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}\big]>\mathbb{E}\big[\|\mathbf{y}-\check{\mathbf{f}}_{k,m,p}\|^{2}\big].$
7 Simulations
To validate Theorems 4.1 and 4.2 and assess their generality beyond orthogonal designs, we conducted a simulation study. We generated IID training data $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$ with $n=1000$, $p=100$, where $\mathbf{x}_{i}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{\Sigma})$ with banded correlation $\Sigma_{ij}=\rho^{|i-j|}$ for $\rho=0.5$, $y_{i}=\boldsymbol{\beta}^{\top}\mathbf{x}_{i}+\varepsilon_{i}$ with $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, and exact sparsity $\beta_{j}=\mathbf{1}(j\leq 10)$. The noise variance $\sigma^{2}$ was chosen to achieve the desired signal-to-noise ratio $\textnormal{SNR}=\boldsymbol{\beta}^{\top}\boldsymbol{\Sigma}\boldsymbol{\beta}/\sigma^{2}$.
For EFS, we used $B=500$ base learners and varied $m$ over a geometric grid from $2$ to $p$, selecting the optimal $m$ via 10-fold cross-validation for each $k\in[20]$. We compared degrees of freedom and training error for EFS($k,m$) against FS($k$) at each model size $k$.
Figure 2 shows results for $\textnormal{SNR}\in\{0.053,0.25,1.0\}$. At low SNR ($0.053$ and $0.25$), EFS($k,m$) reduces both degrees of freedom and training error relative to FS($k$) at each $k$, shifting the entire bias-variance tradeoff curve toward the origin. At high SNR ($1.0$), EFS($k,m$) performs worse near the true sparsity level $k=10$, consistent with Proposition 4.1.
(a) SNR =0.053=0.053
(b) SNR =0.25=0.25
(c) SNR =1.0=1.0
Figure 2: Degrees of freedom vs. training MSE for EFS($k,m$) and FS($k$) with $p/n=0.1$ and banded exact sparsity, as $k$ varies over $\{1,2,\ldots,20\}$ (true sparsity is 10). At each $k$, EFS($k,m$) reduces both quantities, shifting the tradeoff curve toward the origin. Gains are more pronounced at low SNR. At high SNR, EFS($k,m$) and FS($k$) become similar.
8 Discussion
We have revisited the role of feature subsampling in randomized ensembles such as random forests, focusing on subtleties that emerge when subsampling is interleaved with greedy optimization rather than applied to base learners fit via convex optimization. By analyzing ensemble forward selection under orthogonal fixed design, we moved beyond the empirical observations of curth2024random; liu2025randomization; revelas2025random to prove that EFS($k,m$) can simultaneously achieve smaller bias and smaller variance than the base estimator FS($k$). Our work is similar in spirit to mei2024exogenous, who analyzed an oracle version of RFs, but differs in our ability to analyze a non-oracle algorithm, explicitly quantify the performance gap, and characterize the resulting solution. In particular, we reveal a logistic reweighting of the OLS coefficients.